OpenStudy (anonymous):

The ordered pairs (1, 81), (2, 100), (3, 121), (4, 144), and (5, 169) represent a function. What is a rule that represents this function? (1 point)y = x8 can you help me not just give answer

OpenStudy (anonymous):

Look at the pairs. 1 - 81 2 - 100 3 - 121 THe first numbers are just going upwards in succession obviously. 1, 2, 3, 4... But the second numbers are all perfect square. 81 (9x9), 100 (10x10), 121 (11x11). Do you see the pattern?

OpenStudy (anonymous):

plus one

OpenStudy (anonymous):

my options are y = x8 y = 8x y = (x + 8)2 y = 8x

OpenStudy (anonymous):

would it be number 3(x+8)2

OpenStudy (radar):

\[y=(x+8)^{2}\] (1+8)^2 = 81 (2+8)^ = 100 looks like a winner to me.